OpenStudy (coconutjj):

dx/dt = -2 , dy/dt = 3 If z^2=(x^2)(y^2), what is dz/dt when x=2, y=3

jimthompson5910 (jim_thompson5910):

\[\Large z^2 = x^2*y^2\] \[\Large \frac{d}{dt}[z^2] = \frac{d}{dt}[x^2*y^2]\] \[\Large \frac{d}{dt}[z^2] = \frac{d}{dt}[x^2]*y^2+x^2*\frac{d}{dt}[y^2]\] \[\Large 2z*\frac{dz}{dt} = 2x*\frac{dx}{dt}*y^2+x^2*2y*\frac{dy}{dt}\] I'll let you finish up